1. Find the derivative of the given function, f(x)=1966



2. Find the derivative of the following function: f(x)=x²+6x+9



3. Find the slope of the tangent line to the graph of the following function at the indicated point, f(x)=3-2x at (-1, 5)



4. Find the derivative of the function: f(x)=x½



5. Find the derivative of the function f(x)=5x²(x+47)



6.) Find the derivative of the function: f(x)=5(x+47)²



7. Find the derivative of the function f(x)= 5x²/(x+47).



8. Which of the following functions are NOT differentiable?


a. f(x)=lxl


b. f(x)=(x+3)⁴


c. f(x)=mx+b


d. f(x)=1066


e. none of the above



9. Which of the following functions are NOT everywhere continous?


a. f(x)=x²-4/(x+2)


b. f(x)=(x+3)⁴


c. f(x)=1066


d. f(x)=mx+b


e. none of the above




1
Expert's answer
2022-03-11T06:43:16-0500

Q.1:    Find the derivative of the given function, "f(x)=1966"

Solution


"\begin{array}{l}
f\left( x \right) = 1966\\
f'\left( x \right) = 0
\end{array}\"               This is a constant function 




Q.2:    Find the derivative of the following function:"f(x)=x\u00b2+6x+9"

Solution


"\begin{array}{l}
f\left( x \right) = {x^2} + 6x + 9\\
f'\left( x \right) = 2x + 6
\end{array}\"




Q.3:    Find the slope of the tangent line to the graph of the following function at the indicated point, "f(x)=3-2x" at "(-1, 5)"

Solution


"\begin{array}{l}
f\left( x \right) = 3 - 2x\\
f'\left( x \right) = - 2
\end{array}\"


The gradient of the tangent line is "m = f'\left( x \right) = - 2\"


Equation of tangent line is


"\begin{array}{l}
y - {y_1} = m\left( {x - {x_1}} \right)\\
y - 5 = - 2\left( {x + 1} \right)\\
2x + y = 3
\end{array}\"



Q.4:    Find the derivative of the function: "f(x)=x\u00bd"

Solutio


"f\left( x \right) = {x^{{\textstyle{1 \over 2}}}}\"

"f'\left( x \right) = \frac{1}{2}{x^{{\textstyle{1 \over 2}} - 1}}\"

"f'\left( x \right) = \frac{1}{2}{x^{{\textstyle{{ - 1} \over 2}}}}\"



Q.5:    Find the derivative of the function "f(x)=5x\u00b2(x+47)"

Solution


"\begin{array}{l}
f\left( x \right) = 5{x^2}\left( {x + 47} \right)\\
f\left( x \right) = 5{x^3} + 235{x^2}\\
f'\left( x \right) = 15{x^2} + 470x
\end{array}\"



Q.6:    Find the derivative of the function: "f(x)=5(x+47)\u00b2"

Solution


"\begin{array}{l}
f\left( x \right) = 5{\left( {x + 47} \right)^2}\\
f'\left( x \right) = 5 \cdot 2{\left( {x + 47} \right)^{2 - 1}} = 10\left( {x + 47} \right)
\end{array}\"



Q.7:    Find the derivative of the function "f(x)= \frac{5x²}{x+47}"

Solution



"\begin{array}{l}
f\left( x \right) = \frac{{5{x^2}}}{{\left( {x + 47} \right)}}\\
f'\left( x \right) = \frac{{\left( {x + 47} \right){{\left( {5{x^2}} \right)}^\prime } - \left( {5{x^2}} \right){{\left( {x + 47} \right)}^\prime }}}{{{{\left( {x + 47} \right)}^2}}}\\
f'\left( x \right) = \frac{{\left( {x + 47} \right)\left( {10x} \right) - \left( {5{x^2}} \right)\left( 1 \right)}}{{{{\left( {x + 47} \right)}^2}}}\\
f'\left( x \right) = \frac{{10{x^2} + 470x - 5{x^2}}}{{{{\left( {x + 47} \right)}^2}}}\\
f'\left( x \right) = \frac{{5{x^2} + 470x}}{{{{\left( {x + 47} \right)}^2}}}
\end{array}\"


Q.8: Which of the following functions are NOT differentiable?

a. f(x)=lxl

b. f(x)=(x+3)⁴

c. f(x)=mx+b

d. f(x)=1066

e. none of the above

Solution


f(x)=|x| is not differentiable at x = 0 



Q.9:    Which of the following functions are NOT everywhere continuous?

a. f(x)=x²-4/(x+2)

b. f(x)=(x+3)⁴

c. f(x)=1066

d. f(x)=mx+b

e. none of the above

Solution


The correct answer is (e)